OpenStudy (anonymous):

Number of visits, X=x [ 0 ] [ 1 ] [ 2 ] [ 3 ] [ 4 ] Probabllity, P(X=x) [0.1] [0.1] [0.2] [0.4] [0.2] I've figured out the first three parts of the question, but the last part asks, "Given that an adult have fewer than 3 hospital visits, what is the probablity that he or she will have only one hospital visit?" I don't understand what they are asking. The previous answer is .4 or 40% for the probability of having 3 or fewer visits.

OpenStudy (anonymous):

0.1 / (0.1 + 0.1 + 0.2) = 1/4 The main idea behind that Pr (A given B) = Pr (A and B) / Pr B

OpenStudy (anonymous):

Oh, okay. So the probability would be .25 of 25%? The 0.1 of the 1 stay is divided by the 0,1, & 2 stays...correct? I want to understand how o do it :)